OpenStudy (anonymous):

What is the sum of the multiples of 4 from 16 to 100? I got n to be 25 but should I use some other value of n because it starts on the 4th multiple?

OpenStudy (raden):

yeah, the series should be : 16,20,24,...,100 to find the value of n : an = a1+(n-1)d 100 = 16 + (n-1)4 100 = 4n + 12 4n = 100 - 12 4n = 88 n=22

OpenStudy (raden):

now, can u calculate of 16+20+24+28+...+96+100 = ....

OpenStudy (anonymous):

I got 1276

OpenStudy (raden):

11 * (16+100) = 11 * 116 = 1276 u are right